working together, linda and kathy can clean a beach house in 5.14 hours. had she done it alone it would have taken linda 11 hours. how long would it take kathy to do it alone?
How do i set this problem up?

Answer by TimothyLamb(4379)

x = linda work rate
y = kathy work rate
---
rate = work/time
time = work/rate
work = 1 job (clean one house)
---
x = 1/11
x + y = 1/5.14
1/11 + y = 1/5.14
y = 1/5.14 - 1/11
---
answer:
time for kathy working alone = work/rate
1/y = 1/( 1/5.14 - 1/11 )
1/y = 9.65 hours
---
check:
x + y = 1/5.14
1/11 + (1/5.14 - 1/11) = 1/5.14
5.14/(11*5.14) + ( 11/(11*5.14) - 5.14/(11*5.14) ) = 1/5.14
5.14/(11*5.14) + 11/(11*5.14) - 5.14/(11*5.14) = 1/5.14
11/(11*5.14) = 1/5.14
1/5.14 = 1/5.14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php